The perimeter of a rectangle is 50, if the lenght is 5 more than three times it's width, what is the width?
Answer by nerdybill(2710)

The perimeter of a rectangle is 50, if the lenght is 5 more than three times it's width, what is the width?
.
Let w = width
then
3w+5 = length
.
2(w + 3w+5) = 50
(w + 3w+5) = 25
4w + 5 = 25
4w = 20
w = 5 (width)